OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 10, and minor axis of length 8

OpenStudy (michele_laino):

we have to apply this equation: \[\Large \frac{{{x^2}}}{{{a^2}}} + \frac{{{y^2}}}{{{b^2}}} = 1\] where a, and b are such that: \[\Large 2a = 10,\quad 2b = 8\]

OpenStudy (anonymous):

okay